OpenStudy (anonymous):

Solve: 4x − 10 + 16 = 22 I'm getting 12, but my computer keeps getting 4 so I'm not sure what ones right.

OpenStudy (hrhsbull):

x does equal 4 16 divided by 4 is 4

myininaya (myininaya):

Are you doing 16-4 instead of 16/4? I think you are forgetting that to undo multiplication you divide (not divide on one side and subtract on the other ->you can't do this) Whatever you do to one side you must do to the other. If you have 4x=16 You look and see that that x is being multiplied by 4 To undo that multiplication of 4 you divide by 4. And if you divide by 4 on that side, then you do it to the side side. If we have 4+x=16 Then we have addition of 4 and to undo addition by 4 we subtract by 4. If we subtract 4 on one side then we subtract 4 on the other side. So if we have 4x=16, then you do 4x/4=16/4 x = 4 4x/4=1x since 4/4=1 If we have 4+x=16, then you do 4+x-4=16-4 x =12 4+x-4=x since 4-4=0

OpenStudy (anonymous):

Here's how I'm doing it: 4x − 10 + 16 = 22 4x - 26 = 22 4x = 48 48 / 4 = 12 12 = x ----- Substitute 12 for x: 4(12) − 10 + 16 = 22 48 - 10 + 16 = 22 48 - 26 = 22 22 = 22

myininaya (myininaya):

-10+16=6

myininaya (myininaya):

Since we are adding numbers with different signs we subtract. 16-10=6 Since 16>10 and 16 is positive then -10+16=6 OR! you could realize that -10+16=16+(-10)=16-10

OpenStudy (anonymous):

@kathy0514 What do you think too?

myininaya (myininaya):

Do you see that I'm saying that you have been adding wrong?

OpenStudy (anonymous):

Ummm somewhat.... ...not really... OK, so lets start with the equation from the begininng. 4x − 10 + 16 = 22 The first step is to add (+)10 and (+)16 together, right? 10 + 16

myininaya (myininaya):

-10+16=6

myininaya (myininaya):

There is a negative in front of that 10 not a positive.

OpenStudy (anonymous):

4x − 10 + 16 = 22 4x + 6 = 22 4x = 22 - 6 4x = 16 x= 16/4 x= 4

myininaya (myininaya):

Yep very good.

OpenStudy (anonymous):

If you use the - sign with the 10, how do you know what operation to put between 4x and 6?

myininaya (myininaya):

You have subtraction is addition of the opposite 4x+(-10)+16=22

myininaya (myininaya):

4x+(-10+16)=22

myininaya (myininaya):

4x+(6)=22 4x+6=22

OpenStudy (anonymous):

Are you sure? It's supposed to be a "2 step equation".

myininaya (myininaya):

If we had x-2+4=6 Then we can look at it as x+(-2)+4=6 Since (-2) and 4 are like terms I can look at adding them x+(-2+4)=6 x+(2)=6 x+2=6

OpenStudy (anonymous):

Simplifying 4X + 10 = 22 Reorder the terms: 10 + 4X = 22 Solving 10 + 4X = 22 Solving for variable 'X'. Move all terms containing X to the left, all other terms to the right. Add '-10' to each side of the equation. 10 + -10 + 4X = 22 + -10 Combine like terms: 10 + -10 = 0 0 + 4X = 22 + -10 4X = 22 + -10 Combine like terms: 22 + -10 = 12 4X = 12 Divide each side by '4'. X = 3 Simplifying X = 3

OpenStudy (anonymous):

did that help?

myininaya (myininaya):

Were you talking to be? I'm confused. Also I'm totally sure -10+16 is 6.

OpenStudy (anonymous):

@kathy0514 Not really, cause I already know how to do that :P @myininaya @kathy0514 I'm talking to both of you.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

OK, so I have another equation that I'm going to show you guys how I'm doing it. Tell me if I do anything wrong.

myininaya (myininaya):

The equation you gave was 4x-10+16=22 Now this is the same as saying 4x+(-10)+16=22 Now we notice we have like terms, the -10 and the 16.The operation in between -10 and 16 is addition. So we add -10 and 16 giving us the result 6. So -10+16=6 So we can now reduce our equation to this 4x+6=22 Since our objective is to isolate x we will need to undo the addition of 6 by subtracting 6. If we subtract 6 on one side, then we subtract 6 on the other side like this: 4x+6-6=22-6 4x+0 =16 Since 4x+0=4x We have 4x =16 To undo multiplication by 4 we divide by 4.If we divide 4 on one side then we do so to the other side 4x/4=16/4 now since 4/4=1 we have x =16/4 But oh mine 16/4 is 4 so we actually have that x= 4.

OpenStudy (anonymous):

Solve: 4(2x − 1) − 3 = 9(x − 4): 8x - 4 - 3 = 9x - 36 8x + -7 = 9x - 36 17x + -7 = -36 17x = - 29

OpenStudy (anonymous):

4(2x - 1) - 3 = 9(x - 4) 8x -4 -3 = 9x - 36 8x - 7 = - 9x - 36 17x = -29 x = -29/17

myininaya (myininaya):

Well it looks like you subtracted 9x on one side and then added it on the other side. Remember whatever you do to one side do to the other.

OpenStudy (anonymous):

On line 8x - 7 = - 9x - 36 I didn't mean to put the (-) sign in front of the 9.

myininaya (myininaya):

well actually I'm confused. You went from 9x to -9x

myininaya (myininaya):

oh.

myininaya (myininaya):

But still 8x-9x=-x

myininaya (myininaya):

4(2x − 1) − 3 = 9(x − 4) 4(2x)-4(1)+(-3)=9(x)-9(4) 4(2)x-4+(-3)=9x-36 8x - 4 + (-3) = 9x-36 8x - 7 =9x-36 add 7 add 7 8x = 9x-29 subtract 9x subtract 9x -1x = -29 Which means x=29

OpenStudy (anonymous):

4(2x − 1) − 3 = 9(x − 4): 8x - 4 - 3 = 9x - 36 8x (+9) + -7 = 9x (+9) - 36 17x + -7 (- -7) = -36 (- - 7) 17x = - 29

myininaya (myininaya):

9x+9x=18x not 0x

OpenStudy (anonymous):

No, because you add the same thing, it cancels out.

myininaya (myininaya):

No 9x-9x=0

myininaya (myininaya):

You need to subtract 9x on both sides.

myininaya (myininaya):

Or that is one possibility to go about this you could have done the following: 4(2x − 1) − 3 = 9(x − 4) 8x - 4 +(-3) =9x-36 8x + (-7) =9x-36 subtract 8x on both sides -8x -8x (-7) =1x-36 -7 =x -36 Now add 36 on both sides -7 = x -36 +36 +36 29 = x

OpenStudy (anonymous):

@Niall did u understand?

OpenStudy (anonymous):

Sorry, I was doing something. I think I get it. :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do u want a medal?

OpenStudy (anonymous):

Um sure thanks. :) Solve: 4(x + 1) = 4x + 4 No Solution zero Any Real Number 1 For this one I think the answer would be any real number, right?

OpenStudy (anonymous):

Distribute 4 into x + 1, and we have two identical expressions... 4x + 4 = 4x + 4 Noting that both sides are congruent, we have infinite solutions for x. I hope this helps!

OpenStudy (anonymous):

So there is no solution? I'm kind of confused about what is wants the answer to be, because as far as I'm aware, is would just be "True" :P

OpenStudy (anonymous):

u r correct any real number

OpenStudy (anonymous):

thanx 4 the medal

OpenStudy (anonymous):

Ok, thanks! Your Welcome :)

OpenStudy (anonymous):

;o)

OpenStudy (anonymous):

8-)